A spark ignition engine operates on an otto cycle with a compression ratio of 9, and temperature limits of 30 degrees Celsius and 1000 degrees Celsius. If the power input is 500kw,calculate the mass flow rate of air.
1
Expert's answer
2021-07-19T02:52:02-0400

"\eta=1-\frac{1}{9}^{1.4-1}=1-(\frac{1}{9})^{0.4}"

"\eta=1-.41=0.58"

Mass flow rate 0.58